OpenStudy (anonymous):

Convert 1.75 g/mL to the unit g/m3. 1.75 x 106 g/m3 1.75 x 102 g/m3 1.75 x 10-2 g/m3 1.75 x 10-6 g/m3

OpenStudy (anonymous):

I really do not understand anything about this problem besides that you have to cube the conversion. I am unsure of what the conversion is because g/mL and g/m^3 look like different measurement types completely. Liter -> Meter

OpenStudy (anonymous):

One milliliter equals one cubic centimeter. Since you're going from cubic centimeters to cubic meters, you need to change the length units by a factor of 100. You're correct in recognizing that you need to cube the conversion factor. In this case, your dimensional analysis looks like this: \[\frac{ 1.75g }{ 1 mL } * \frac{ 1mL }{ 1cm^3 }*\frac{ 100^3cm^3 }{ m^3 }=1.75*10^6\]The key thing to remember here is the conversion from milliliters to cubic centimeters. From there, it's a relatively straightforward change-of-units problem.